# LACONIC: A 3D Layout Adapter for Controllable Image Creation

### v1

@@ 2 Related Work
”A square bedroom with a king-size bed, flanked by two nightstands. A wardrobe, approximately the same width as the bed, stands in front of it, with decorative items inside. Next to one of the nightstands, there is a large Alocasia plant. A window is positioned in front of the plant and beside the wardrobe. The plant is at the forefront, partially obstructing the view.”
@@ 4 Experiments
We follow previous work [73] and use HyperSim [52] indoor scene dataset to extract semantic 3D bounding box layouts as well as camera poses with their associated photorealistic renderings. As HyperSim only features $326$ unique 3D layouts, paired to a total of $24{,}383$ images, methods taking as input the comprehensive 3D layout associated to each image sample, such as ours, may tend to memorize individual fixed scenes from the dataset, making it difficult to perform the complex out-of-distribution manipulations of individual objects that our framework enables. In response, and in order to provide additional qualitative results showcasing the full range of our model’s capabilities, we also gather a custom dataset featuring $72{,}000$ bedroom 3D scene layouts, each being paired to a single, high-quality image and viewpoint. See supplementary for additional details and statistics on the used datasets.
@@
Figure 5: Impact of the adapter strength on generated images. From a text caption $c=$ ”A Retro Synthwave Indoor Bedroom”, adherence with the input 3D layout improves accordingly with the parameter $\gamma$ from Equation (3), achieving strict control at higher scales. We can observe that even at lower scales (c) the control enforces the geometric and semantic plausibility of the synthesized image.
Throughout our experiments, we employ Stable Diffusion v1.5 [53] as the pretrained text-to-image diffusion backbone upon which our adapter network is trained. We train the DM-FS baseline using a similar but downsized UNet architecture, to account for the scale of the dataset. SceneCraft [73] is used with its pretrained weights and default parameters from the official implementation. All the compared methods in quantitative evaluations and side-by-side qualitative comparisons are trained on the same HyperSim [52] subset. We provide comprehensive implementation details as well as training and inference settings in the supplementary materials.
@@
Figure 6: Illustration of our metric’s object selection mechanism. Obstructed and out-of-bounds objects are identified from their 2D and 3D bounding box annotations and filtered out from the evaluation set.
@@
Since each visible object $o_{i}$ is associated with a text description, we can compute a semantic correlation between the object’s text caption $s_{i}$ and its crop image $x_{o_{i}}$. For this last step, we follow [15] among others, and use the pretrained CLIP [51] model. We provide details on our Scene Object Clip score (SOC) in the supplementary materials.
@@ Results
Our main quantitative results are summarized in Table 1. We report results both with and without text prompting. In the latter case our approach achieves state-of-the-art results across all metrics and outperforms the very recent baseline SceneCraft [73] by a significant margin. When using text conditioning (the captions provided in HyperSim [52]), our method still outperforms SceneCraft [73] on most metrics and in several cases by a significant margin. We note that our approach is able to capture the HyperSim data distribution significantly better compared to SceneCraft [73]. However, due to the relative limited scope of 3D structures in that dataset, our approach demonstrates somewhat less variety in text-driven scenarios as indicated by a lower Inception Score. Importantly, our adapter-based approach outperforms the DM-FS approach on all metrics, while enabling text-conditioned synthesis, thus validating the importance of using a rich pretrained T2I model. We provide qualitative comparisons on the 3D layout-guided image synthesis task in Figure 4. Notably, we observe that our approach produces more natural images, featuring fewer visual artifacts compared to baseline methods, while being more faithful to the input 3D layout and viewpoint.
@@ 4.2 Additional Qualitative Results
We provide additional generation results, showcasing the intuitive and iterative editing capabilities of our framework as well as its ability to leverage the pretrained text-to-image backbone for out-of-distribution generalization in Figure LABEL:fig:teaser and Figure 3. Overall, these results highlight that our model can capture nuanced individual object relations within structured scenes. Moreover, our approach enables manipulations in 3D space, and provides wholistic scene modeling, accounting, e.g., for off-screen objects. Finally, results reported in Figure 5 show how introducing our adapter control, even at low scale $\gamma$, allows synthesizing 3D-aware images that are more geometrically and semantically sound than base text-to-image outputs, while maintaining adherence with the textual description. Additional qualitative results highlighting the power of our approach can be found in the supplementary materials.
@@ 5 Conclusion, Limitations & Future Work
Figure 7: Additional image synthesis results (Right) from an input 3D layout and prompt (Left). Our method shows great adherence to both input 3D layout and caption prompt. Notably, LACONIC generates high quality results without needing prompt information (bottom center), with in-domain prompts (top right & bottom left) and with out-of-distribution prompts (top left/center & bottom right).
@@ Appendix A Implementation
In this section, we provide comprehensive implementation details on our 3D layout adapter architecture, pretrained text-to-image diffusion backbone, as well as our training and inference settings. Unless explicitly specified, same settings are shared across the different datasets that we employed to conduct our experiments. We also detail the training and test configurations of the baseline methods.
@@ A.1.1 Pretrained Models
We employ the established and widely adopted Stable Diffusion [53] v1.5 to implement the conditional UNet image denoiser $\epsilon_{\theta}$ and CLIP [51] text prompt encoder $\tau_{\theta}$, using the pretrained weights and implementation [^1] https://huggingface.co/stable-diffusion-v1-5/stable-diffusion-v1-5 from the HuggingFace Diffusers [66] library. In our framework, the text encoder is used to encode both the global image caption $c$ and the object-level semantic descriptions $s$.
@@
We detail the individual trainable modules that, building upon previous work [70, 36, 44], are used to embed the input semantic 3D layout $\mathcal{S}$, as illustrated in Figure 2 and described in Section 3.3 in the main paper.
@@
Parametric object bounding box representation $o_{i}$ are embedded by a common module. More precisely, each scalar defining individual object spatial attributes $(p_{i},R_{i},d_{i})$ is projected to a higher dimensional vector using fixed, sinusoidal positional encoding [65] with $32$ frequencies. We follow [70] and apply:
@@
Consequently, object’s position $p_{i}$ and dimension $d_{i}$ are encoded to $192$-dimensional attributes. Importantly, the rotation matrix $R_{i}$ is first expressed in a continuous representation, following the recommendations from [81], leading to a vector in $\mathbb{R}^{6}$ from which individual scalars are also processed by Equation (5). This produces a representation in $\mathbb{R}^{384}$ that is subsequently mapped by a linear layer in $\mathbb{R}^{192}$. The semantic embedding $\tau(s_{i})$ is the end-of-sequence output token from the pretrained text encoder in $\mathbb{R}^{d_{\tau}}$, with $d_{\tau}=768$, and is further projected to a $192$-dimensional representation by an MLP featuring a single hidden unit of dimension $384$ with LeakyReLU activations. The mentioned encoded object attributes are finally concatenated, yielding an individual token $\mathcal{T}_{o_{i}}$ of dimension $4\times 192=d_{\tau}$ for each of the $N$ objects in the scene.
@@
Following prior work [70, 36], we encode the optional scene’s floor plan $\mathcal{F}$ by leveraging a popular implementation [^2] https://github.com/fxia22/pointnet.pytorch of the PointNet [50] model, applied on a point cloud representation obtained by sampling $P=100$ three-dimensional points, that are evenly spaced along the room’s boundaries. A single floor token $\mathcal{T}_{\mathcal{F}_{i}}\in\mathbb{R}^{d_{\tau}}$ is subsequently obtained by projecting the $1024$-dimensional module’s output.
@@
New token representations $\hat{\mathcal{T}}$ are established by a Transformer encoder that follows the seminal paper [65] by using the implementation from the PyTorch [45] library. In practice, we use $4$ encoder layers with $6$ attention heads and hidden dimensions of size $512$. Since the scene is defined as an unordered sequence of its objects and floor representations, we do not additionally encode the position of individual tokens in $\mathcal{T}$, and perform zero-padding to handle scenes with different number of objects $N$. In practice, we consider a maximum number of $50$ objects. Following [75], the output token embeddings are each passed to a shared MLP that preserves the token dimension $d_{\tau}$, implemented with a hidden unit of size $768$, GELU activation [14], and followed by Layer Normalization [1].
@@ A.1.3 Cross-Attention Layers
As introduced in Section 3.3 in the main paper, the scene conditioning sequence $\hat{\mathcal{T}}$ outputted by the 3D layout encoder is mapped to associated key $K^{y}$ and value $V^{y}$ by introducing respective learnable dense projection matrices $W_{K}\in\mathbb{R}^{d\times d_{\tau}}$ and $W_{V}\in\mathbb{R}^{d\times d_{\tau}}$. In practice, we follow the implementation [^3] https://github.com/tencent-ailab/IP-Adapter from IP-Adapter [75] to introduce the decoupled cross-attention mechanism within residual blocks from the pretrained Stable Diffusion UNet backbone.
@@ A.2 Datasets
In this section, we introduce the datasets that we used to conduct our experimental evaluations and introduce any relevant dataset-dependent mechanisms implemented to adapt our work to their specific features or available annotations.
@@ A.2.1 HyperSim Dataset
Following previous work [73], we leverage indoor scenes and photorealistic renderings from HyperSim [52] to construct a collection of semantic 3D bounding box layouts with camera poses $y$ and image $x$ pairs. The dataset originally features $461$ scenes from diverse types of indoor environments, and $77{,}400$ rendered images. We follow SceneCraft [73] and leverage a filtered version [^4] https://huggingface.co/datasets/gzzyyxy/layout_diffusion_hypersim of the dataset proposed in [58], discarding unbounded and scenes with excessively large scales. This results in $323$ unique scenes, that are associated to $24{,}383$ images annotated by a global text caption extracted by a pretrained foundation model. Each layout features 3D bounding box annotations that are typed according to the standard NYU40 [60] semantic classes. Each rendering is also associated to an object instance map, that we leverage in the context of our SOC metric implementation, as detailed in Section A.5. We conducted foundational statistical analysis of the 3D layouts, establishing that scenes contain an average of $121$ objects, with a median of $54$ and that the dataset demonstrates a large diversity in terms of structural complexity, which is challenging in the context of our work and in light of the limited number of available samples.
@@
We describe here the dataset-specific mechanisms that we implemented to run experiments on HyperSim layouts. Importantly, a proportion of scenes from the dataset feature hundreds of objects, exceeding the maximum length fixed to $50$ for our sequence-based representation used for the model conditioning. As a result, we leverage the instance semantic map of the target rendering to prioritize including visible objects in the conditioning sequence. In case the number of visible objects still exceeds the maximum length, we identify the most meaningful objects based on their bounding box dimensions. As the HyperSim dataset does not provide the floor information in a straightforward manner, we set $\mathcal{F}=\emptyset$ in our experiments. Object-level semantic information are also limited to their categorical label among NYU40 classes, from which we apply $\tau_{\theta}(s_{i})$ as described in our general methodology.
@@ A.2.2 Custom Bedroom Dataset
As described in the main paper in Section 4.1, the limited number of available HyperSim layouts, that are passed as conditioning input, makes it easier for our model to learn a mapping between the global scenes and specific camera poses and target renderings, while overlooking the contributions of individual objects. In response and for the purpose of our qualitative evaluations, we gathered a custom dataset of $76{,}000$ human-designed 3D indoor bedroom layouts, in which each layout is associated to a single rendering. The dataset also includes floor information, represented as a 3D point cloud. Finally, ground-truth object-level semantic descriptions $s_{i}$ are extracted using a LLaVA model [29] from their 2D rendering using the following instruction: ”Describe this object concisely. Do not analyse the background, only the object”. This level of annotations allows to implement all the building blocks from our general methodology.
@@ A.3 Training Protocol
To optimize our adapter network following a diffusion-based approach with frozen text-to-image denoiser, we follow standard DDPM [18] training with $\epsilon$-prediction objective, as formalized by Equation (1) in the main paper, and uniformly sampled noise timesteps $t\sim\mathcal{U}(0,1000)$. We randomly drop the conditioning input 3D layout $y$ with a rate $\mathbf{p}_{\text{drop}}=0.15$. We use the AdamW [32] optimizer, with weight decay coefficient $\lambda=0.01$ and learning rate $\eta=5\times 10^{-5}$, reached following an initial linear warmup phase during the first $200$ training iterations. Our checkpoints are trained for a total of $250$ epochs, with $\eta$ decreasing according to a cosine schedule. We leverage PyTorch Lightning [8] to distribute experiments across $3$ NVIDIA RTX 6000 GPUs, each handling a batch size of $10$ samples. Importantly, we experimented enabling automatic mixed precision during training, but empirically observed instability and inconsistent convergence.
A.4 Baselines
A.4.1 SceneCraft
SceneCraft [73] is a recent baseline method that proposes to tackle 3D layout-guided image synthesis, in which, similar to our approach, 3D layouts are defined as a collection of typed object 3D bounding boxes. To do so, it renders the input 3D layout from the target camera viewpoint to (i) a semantic map in which individual boxes are colored according to their one-hot semantic category and (ii) depth maps that are automatically derived from the input geometry. From this pair of 2D input representations, SceneCraft trains dedicated ControlNet [78] modules in supervised experiments and in order to introduce the additional controls to a pretrained text-to-image backbone. In practice, we leverage the pretrained checkpoint [^5] https://huggingface.co/gzzyyxy/layout_diffusion_hypersim_prompt_one_hot_multi_control_bs32_epoch24 obtained by training the modules on the same HyperSim [52] subset that we used in our own experiments. It uses Stable Diffusion [53] v2.1 as the T2I prior. We employ official author implementation [^6] https://github.com/OrangeSodahub/SceneCraft with the default test-time hyperparameters.
A.4.2 Diffusion Model from Scratch
We describe below the main motivations and implementation details behind the Diffusion Model trained From Stracth (DM-FS) from our experimental evaluations.

### v2

@@ 2 Related Work
“A square bedroom with a king-size bed, flanked by two nightstands. A wardrobe, approximately the same width as the bed, stands in front of it, with decorative items inside. Next to one of the nightstands, there is a large Alocasia plant. A window is positioned in front of the plant and beside the wardrobe. The plant is at the forefront, partially obstructing the view.”
@@ 4 Experiments
We follow previous work [77] and use HyperSim [54] indoor scene dataset to extract semantic 3D bounding box layouts as well as camera poses with their associated photorealistic renderings. As HyperSim only features 326326326 unique 3D layouts, paired to a total of 24,38324{,}38324 , 383 images, methods taking as input the comprehensive 3D layout associated to each image sample, such as ours, may tend to memorize individual fixed scenes from the dataset, making it difficult to perform the complex out-of-distribution manipulations of individual objects that our framework enables. In response, and in order to provide additional qualitative results showcasing the full range of our model’s capabilities, we also gather a custom dataset featuring 72,00072{,}00072 , 000 bedroom 3D scene layouts, each being paired to a single, high-quality image and viewpoint. See the Supplementary Material Section A.2 for additional details and statistics on the used datasets.
@@
Figure 5: Impact of the adapter strength on generated images. From a text caption $c=$ “A Retro Synthwave Indoor Bedroom”, adherence with the input 3D layout improves accordingly with the parameter $\gamma$ from Equation (3), achieving strict control at higher scales. We can observe that even at lower scales (c) the control enforces the geometric and semantic plausibility of the synthesized image.
Throughout our experiments, we employ Stable Diffusion v1.5 [55] as the pretrained text-to-image diffusion backbone upon which our adapter network is trained. We train the DM-FS baseline using a similar but downsized UNet architecture, to account for the scale of the dataset. SceneCraft [77] is used with its pretrained weights and default parameters from the official implementation. All the compared methods in quantitative evaluations and side-by-side qualitative comparisons are trained on the same HyperSim [54] subset. We provide comprehensive implementation details, training and inference settings, and an ablation study in the supplementary material.
@@
Figure 6: Object selection for the SOC metric. Obstructed and out-of-bounds objects are identified from their 2D and 3D bounding box annotations and filtered out from the evaluation set.
@@
Since each visible object $o_{i}$ is associated with a text description, we can compute a semantic correlation between the object’s text caption $s_{i}$ and its crop image $x_{o_{i}}$. For this last step, we follow [16] among others, and use the pretrained CLIP [53] model. We provide details on our Scene Object Clip score (SOC) in the supplementary Section A.6.
@@ Results
Our main quantitative results are summarized in Table 1. We report results both with and without text prompting. In the latter case our approach achieves state-of-the-art results across all metrics and outperforms the very recent baseline SceneCraft [77] by a significant margin. When using text conditioning (the captions provided in HyperSim [54]), our method still outperforms SceneCraft [77] on most metrics and in several cases by a significant margin. We note that our approach is able to capture the HyperSim data distribution significantly better compared to SceneCraft [77]. However, due to the relative limited scope of 3D structures in that dataset, our approach demonstrates somewhat less variety in text-driven scenarios as indicated by a lower Inception Score. Importantly, our adapter-based approach outperforms the DM-FS approach on all metrics, while enabling text-conditioned synthesis, thus validating the importance of using a rich pretrained T2I model. We provide qualitative comparisons on the 3D layout-guided image synthesis task in Figure 4. Notably, we observe that our approach produces more natural images, featuring fewer visual artifacts compared to baseline methods, while being more faithful to the input 3D layout and viewpoint. These qualitative findings are confirmed by a perceptual study detailed in the Supplementary Material Section C.3.
@@ 4.2 Additional Qualitative Results
We provide additional generation results, showcasing the intuitive and iterative editing capabilities of our framework as well as its ability to leverage the pretrained text-to-image backbone for out-of-distribution generalization in Figure LABEL:fig:teaser and Figure 3. Overall, these results highlight that our model can capture nuanced individual object relations within structured scenes. Moreover, our approach enables manipulations in 3D space, and provides wholistic scene modeling, accounting, e.g., for off-screen objects. Finally, results reported in Figure 5 show how introducing our adapter control, even at low scale $\gamma$, allows synthesizing 3D-aware images that are more geometrically and semantically sound than base text-to-image outputs, while maintaining adherence with the textual description. Additional qualitative results can be found in the supplementary, including a comparison against Build-A-Scene [7] in Section C.2 and, most notably, a demonstration of our adapter’s compatibility with a modern DiT backbone [8] in Section B.
@@ 5 Conclusion, Limitations & Future Work
Acknowledgments
We thank the anonymous reviewers for their insights and suggestions. We also thank Farah Ellouze and Ana Marcusanu for their feedback on the paper draft. This work was supported by Dassault Systèmes SE. The views and conclusions contained in the paper are those of the authors and should not be interpreted as representing official policies, either expressed or implied, of the company.
@@ Acknowledgments
Figure 7: Additional layout-guided image synthesis results from diverse prompts. The generated images (Right) from the input 3D layouts (Left) and text prompts demonstrate our method’s strong adherence to both conditions. Notably, LACONIC produces high-quality results across various settings: without a text prompt (c), with in-domain prompts (e, f), and with out-of-distribution prompts (a, b, d).
@@ Appendix A Implementation
In this section, we provide comprehensive implementation details of our 3D layout adapter architecture, pretrained text-to-image diffusion backbone, and our training and inference settings. Unless otherwise specified, the same settings are used for all datasets in our experiments. We also detail the training and test configurations of the baselines.
@@ A.1.1 Pretrained Models
We employ the established and widely adopted Stable Diffusion [55] v1.5 to implement the conditional UNet image denoiser $\epsilon_{\theta}$ and CLIP [53] text prompt encoder $\tau_{\theta}$, using the pretrained weights and implementation from the HuggingFace Diffusers [70] library. In our framework, the text encoder is used to encode both the global image caption $c$ and the object-level semantic descriptions $s$.
@@
We detail the modules used to embed the input semantic 3D layout $\mathcal{S}$, which build upon previous work [74, 38, 46] and are illustrated in Figure 2 and described in Section 3.3 in the main paper.
@@
Parametric object bounding box representations $o_{i}$ are embedded by a common module. More precisely, each scalar defining individual object spatial attributes $(p_{i},R_{i},d_{i})$ is projected to a higher-dimensional vector using fixed, sinusoidal positional encoding [69] with 323232 frequencies. We follow [74] and apply:
@@
Consequently, object’s position $p_{i}$ and dimension $d_{i}$ are encoded to 192192192-dimensional attributes. Importantly, the rotation matrix $R_{i}$ is first expressed in a continuous representation, following the recommendations from [85], leading to a vector in $\mathbb{R}^{6}$ from which individual scalars are also processed by Equation (5). This produces a representation in $\mathbb{R}^{384}$ that is subsequently mapped by a linear layer to $\mathbb{R}^{192}$. The semantic embedding $\tau(s_{i})$ is the end-of-sequence output token from the pretrained text encoder in $\mathbb{R}^{d_{\tau}}$, with $d_{\tau}=768$, and is further projected to a 192192192-dimensional representation by an MLP featuring a single hidden layer of dimension 384384384 with LeakyReLU activations. Encoded object attributes are finally concatenated, yielding an individual token $\mathcal{T}_{o_{i}}$ of dimension $4\times 192=d_{\tau}$ for each of the $N$ objects in the scene.
@@
Following prior work [74, 38], we encode the optional scene’s floor plan $\mathcal{F}$ by leveraging a popular implementation of the PointNet [52] model, applied on a point cloud representation obtained by sampling $P=100$ three-dimensional points that are evenly spaced along the room’s boundaries. A single floor token $\mathcal{T}_{\mathcal{F}_{i}}\in\mathbb{R}^{d_{\tau}}$ is subsequently obtained by projecting the 102410241024-dimensional module’s output.
@@
New token representations $\hat{\mathcal{T}}$ are established by a Transformer encoder that follows the seminal paper [69] by using the implementation from the PyTorch [47] library. In practice, we use 444 encoder layers with 666 attention heads and hidden dimensions of size 512512512. Since the scene is defined as an unordered sequence of its objects and floor representations, we do not additionally encode the position of individual tokens in $\mathcal{T}$, and perform zero-padding to handle scenes with different number of objects $N$. In practice, we consider a maximum number of 505050 objects. Following [79], the output token embeddings are each passed to a shared MLP that preserves the token dimension $d_{\tau}$, implemented with a hidden unit of size 768768768, GELU activation [15], and followed by Layer Normalization [1].
@@ A.1.3 Cross-Attention Layers
As introduced in Section 3.3 of the main paper, the scene conditioning sequence $\hat{\mathcal{T}}$ output by the 3D layout encoder is mapped to associated key $K^{y}$ and value $V^{y}$ by introducing respective learnable dense projection matrices $W_{K}\in\mathbb{R}^{d\times d_{\tau}}$ and $W_{V}\in\mathbb{R}^{d\times d_{\tau}}$. In practice, we follow the implementation from IP-Adapter [79] to introduce the decoupled cross-attention mechanism within residual blocks of the pretrained Stable Diffusion UNet backbone.
@@ A.2 Datasets
In this section, we introduce the datasets used in our experimental evaluation and describe any dataset-dependent mechanisms implemented to adapt our work to their specific features or available annotations.
@@ A.2.1 HyperSim Dataset
Following previous work [77], we leverage indoor scenes and photorealistic renderings from HyperSim [54] to construct a collection of semantic 3D bounding box layouts with camera poses $y$ and image $x$ pairs. The dataset originally features 461461461 scenes from diverse types of indoor environments, and 77,40077{,}40077 , 400 rendered images. We follow SceneCraft [77] and leverage a filtered version of the dataset proposed in [61], discarding unbounded and scenes with excessively large scales. This results in 323323323 unique scenes, that are associated with 24,38324{,}38324 , 383 images annotated by a global text caption extracted by a pretrained foundation model. Each layout features 3D bounding box annotations that are typed according to the standard NYU40 [63] semantic classes. Each rendering is also associated with an object instance map, that we leverage in the context of our SOC metric implementation, as detailed in Section A.6. We conducted foundational statistical analysis of the 3D layouts, establishing that scenes contain an average of 121121121 objects, with a median of 545454 and that the dataset demonstrates a large diversity in terms of structural complexity, which is challenging in the context of our work and in light of the limited number of available samples.
@@
We describe here the dataset-specific mechanisms implemented to run experiments on HyperSim layouts. Importantly, a proportion of scenes from the dataset contain hundreds of objects, exceeding the maximum length fixed to 505050 for our sequence-based representation used for the model conditioning. As a result, we leverage the instance semantic map of the target rendering to prioritize including visible objects in the conditioning sequence. If the number of visible objects still exceeds the limit, we identify the most meaningful objects based on their bounding box dimensions. Since the HyperSim dataset does not provide the floor information in a straightforward manner, we set $\mathcal{F}=\emptyset$ in our experiments. Object-level semantic information are also limited to their categorical label among NYU40 classes, from which we apply $\tau_{\theta}(s_{i})$ as described in our general methodology.
@@ A.2.2 Custom Bedroom Dataset
As described in the main paper in Section 4.1, the limited number of available HyperSim layouts, that are passed as conditioning input, makes it easier for our model to learn a mapping between the global scenes and specific camera poses and target renderings, while overlooking the contributions of individual objects. In response, we gathered a custom dataset of 72,00072{,}00072 , 000 human-designed 3D indoor bedroom layouts, in which each layout is associated to a single rendering. The dataset also includes floor information, represented as a 3D point cloud. Finally, ground-truth object-level semantic descriptions $s_{i}$ are extracted using a LLaVA model [30] from their 2D rendering using the following instruction: “Describe this object concisely. Do not analyse the background, only the object”. This level of annotations allows to implement all the building blocks from our general methodology.
@@ A.3 Training Protocol
We optimize our adapter network while keeping the text-to-image denoiser frozen by following standard DDPM [19] training with $\epsilon$-prediction objective, as formalized in Equation (1) in the main paper. Noise timesteps are uniformly sampled from $t\sim\mathcal{U}(0,1000)$. At each iteration, we randomly drop the conditioning input 3D layout $y$ with a rate $\mathbf{p}_{\text{drop}}=0.15$. We use the AdamW [33] optimizer, with weight decay coefficient $\lambda=0.01$ and learning rate $\eta=5\times 10^{-5}$, reached following an initial linear warmup phase during the first 200200200 training iterations. Our checkpoints are trained for a total of 250250250 epochs, with $\eta$ decreasing according to a cosine schedule. We leverage PyTorch Lightning [9] to distribute experiments across 333 NVIDIA RTX 6000 GPUs, each handling a batch size of 101010 samples. Importantly, we experimented with enabling automatic mixed precision during training, but empirically observed instability and inconsistent convergence.
A.4 Inference Settings
At test-time, images for quantitative evaluation are generated following the sampling algorithm from DPM-Solver [34]. We apply 303030 denoising steps with a classifier-free guidance scale of 5.05.05.0. Images in our qualitative results are obtained with the DDIM sampler [65] using 505050 steps.
A.5 Baselines
Figure 8: Text-driven DiT synthesis results. Given an input 3D layout, viewpoint and caption (a), LACONIC with Stable Diffusion 3 [8] supports adjusting the adapter strength to balance fidelity to the text prompt against adherence to the input layout (b–d).
A.5.1 SceneCraft
SceneCraft [77] is a recent baseline method that proposes to tackle 3D layout-guided image synthesis, in which, similar to our approach, 3D layouts are defined as a collection of typed object 3D bounding boxes. To do so, it renders the input 3D layout from the target camera viewpoint to (i) a semantic map in which individual boxes are colored according to their one-hot semantic category and (ii) depth maps that are automatically derived from the input geometry. From this pair of 2D input representations and in order to introduce the additional controls to a pretrained text-to-image backbone, SceneCraft trains dedicated ControlNet [82] modules in a supervised experiment. In practice, we leverage the pretrained checkpoint obtained by training the modules on the same HyperSim [54] subset used in our own experiments. It uses Stable Diffusion [55] v2.1 as the T2I prior. We employ official author implementation with the default test-time parameters.
A.5.2 Diffusion Model from Scratch
We describe below the main motivations and implementation details behind the Diffusion Model trained From Scratch (DM-FS) from our experimental evaluations.